Two diamonds free fall from rest from same heights one second apart .How long after the first diamond begins to fall will the two diamond be ten m apart?

Grade:12

1 Answers

somi teez
6 years ago
hi,
from basic equation we can approach this question,
S1 -S2=1/2 gt^2 -1/2 g (t-1)^2
           =1/2 g(2t -1)
a/q,
10 m=1/2 *10 *(2t -1)
2 t= 3 sec
so, t= 1.5 sec
plz approve it ….feel good to discuss more....
best of luck!!